OpenStudy (anonymous):

Jorge's rectangular garden was 17 feet long and had a perimeter of 54 feet. He made his garden 5 feet longer. What is the area of his new garden? A. 22 square feet B. 64 square feet C. 85 square feet D. 220 square feet

OpenStudy (anonymous):

b??

OpenStudy (anonymous):

@J-bird @woooooooooowwwww

OpenStudy (anonymous):

help

OpenStudy (anonymous):

original width=54/2-17=10

OpenStudy (anonymous):

D

OpenStudy (anonymous):

new area=(17+5)*10=220

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

220

OpenStudy (anonymous):

d ok

OpenStudy (anonymous):

 This may help.